OpenStudy (jennychan12):

dy/dx of -1/a (ln((a+bx)/x) ?

OpenStudy (jennychan12):

find dy/dx of \[y = -\frac{ 1 }{ a }\ln (\frac{ a+bx }{ x })\]

OpenStudy (jennychan12):

a and b are constants

OpenStudy (jennychan12):

\[\frac{ dy }{ dx } = -\frac{ 1 }{ a }(\frac{ x }{ a+bx })(\frac{ bx-a-bx }{ x^2 })\] is that right?

sam (.sam.):

Nope, maybe if you can simplify more then the answer is \[\frac{1}{a x+b x^2}\]

OpenStudy (jennychan12):

i have the answer, i just need the work... so my work was incorrect?

sam (.sam.):

From product rule you get \[\frac{dy}{dx}=-\frac{1}{a}(-\frac{a}{ax+bx^2})+\ln(\frac{a+bx}{x})\frac{1}{a^2}\]

OpenStudy (jennychan12):

ohhhhh i just figured it out. nvm i think i got it.

sam (.sam.):

There should be 2 terms

OpenStudy (jennychan12):

yeah i kinda figured after you mentioned it.